OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 3x

zepdrix (zepdrix):

\[\large \cos(3x)=\cos(2x+x)\]From here we can apply the Angle Addition Formula for Cosine.\[\large \cos(2x+x)=\color{cornflowerblue}{\cos(2x)}\cos x-\color{orangered}{\sin(2x)} \sin x\]And from here, we can apply the Cosine Double Angle Formula to the \(\color{cornflowerblue}{\cos(2x)}\) and the Sine Double Angle Formula to the \(\color{orangered}{\sin(2x)}\).\[\large =\left(\color{cornflowerblue}{\cos^2x-\sin^2x}\right)\cos x-\left(\color{orangered}{2\sin x \cos x}\right)\sin x\] Andddd you can simplify it down a bit from here. But everything is in terms of an Angle X instead of 2x or 3x now, so that's good!

zepdrix (zepdrix):

Oh that's really annoying... They said rewrite with ONLY sinx and cosx.... yet all of your multiple choices have sin2x or sin3x in them... what a stupid question :\ Hmmmm

zepdrix (zepdrix):

Oh maybe I'm reading it incorrectly because powers don't copy paste very well. Like option number 1, is that suppose to be: \[\large \cos x-4\cos x \sin^2x\] Or\[\large \cos x-4\cos x \sin (2x)\]

zepdrix (zepdrix):

I was asking a question....